OpenStudy (anonymous):

A rectangular prism has a volume of 6x^4 - 13x^3 - 5x^2. What expressions can represent the dimensions of the prism? Use factoring

hero (hero):

Hint: Factor out x^2 6x^4 - 13x^3 - 5x^2 = x^2(6x^2 - 13x - 5) Then factor 6x^2 - 13x - 5

OpenStudy (anonymous):

x^2, 3x+1, 2x-5 are the answers but my word am I ever clueless on how it got there

hero (hero):

If you knew how to factor, you would know that 6x^2 - 13x - 5 = (3x + 1)(2x - 5) and that 6x^4 - 13x^3 - 5x^2 = x^2(6x^2 - 13x - 5) = x^2(6x^2 - (15 -2)x - 5) = x^2(6x^2 - 15x + 2x - 5) = x^2(3x(2x - 5) + 1(2x - 5)) = x^2(2x - 5)(3x + 1)

OpenStudy (anonymous):

Ohhhh my goodness I get it now it made no sense at first you are a freaking rockstar!~!!!! Thank you so much!!!!!

hero (hero):

What is it that you understand about it now that you didn't before?